find the rth term of an ap .,sum of whose first terms is 2n+3n*n

Grade:11

1 Answers

Arun
6 years ago
 
As we know
Tn = Sn - Sn-1
Similarily
Tr = Sr - Sr-1
Hence
Tr = 2r + 3r^2 - (2(r-1) + 3(r-1)^2)
Tr = 2r + 3r^2 -2r +2 - 3r^2 -3 + 6r
Tr = -1 +6r